Solve ( D3 – 3DD′2+2D′3 )Z = e2x – y


1
Expert's answer
2022-01-18T16:29:27-0500

"( D^3 \u2013 3DD'^2+2D'^3)Z = e^{2x \u2013 y}"

Replace D with m and D' with 1

Auxiliary equation becomes "m^3-3m+2=0"

"(m-1)(m^2+m-2)=0"

(m-1)(m-1)(m+2)=0

m=1,1,-2

"C.F=f_1(y+x)+xf_2(y+x)+f_3(y-2x)"

For particular integral;

"P.I=\frac{1}{( D^3 – 3DD'^2+2D'^3)} e^{2x – y}"

Put D=2 and D'=-1

"P.I=\frac{1}{( 2^3 – 3(2)(-1)^2+2(-1)^3)} e^{2x – y}"

"=\frac{1}{8-6-2)}e^{2x-y}"

We then differentiate the denominator w.r.t D and multiply by x to get,

"P.I=\frac{x}{\frac{d}{dD}( D^3 – 3DD'^2+2D'^3)} e^{2x – y}"

"=\frac{x}{( 3D^2 – 3D'^2+0)} e^{2x – y}"

"=x\frac{1}{( 3D^2 – 3D'^2)} e^{2x – y}"

Now put D=2 and D'=-1

"=x\frac{1}{( 3(2)^2 – 3(-1)^2)} e^{2x – y}\newline=x\frac{1}{( 12– 3)} e^{2x – y}"

"P.I=\frac{x}{9}e^{2x-y}"

The solution to the equation becomes,

z=C.F+P.I

"=f_1(y+x)+xf_2(y+x)+f_3(y-2x)+\frac{1}{9}xe^{2x-y}"